OpenStudy (anonymous):

**Calc 2 Help Needed** Solve the initial value problem below: (t^2 + 2t)dx/dt = 2x + 5 with t, x>0 and x(1) = 1

OpenStudy (dumbcow):

rearrange terms to separate the x variables from t variables \[\frac{dx}{2x+5} = \frac{dt}{t^{2}+2t}\] integrate both sides \[\frac{1}{2}\ln (2x+5) = \frac{1}{2}(\ln t -\ln (t+2) +C)\] Let C = ln(k) , combine logs \[\ln(2x+5) = \ln(\frac{kt}{t+2})\] \[2x+5 = \frac{kt}{t+2}\] \[x = \frac{kt}{2(t+2)}-\frac{5}{2}\] solve for k using x(1)=1 \[1 = \frac{k}{6}-\frac{5}{2}\] \[k = 21\] \[x(t) = \frac{21t}{2(t+2)}-\frac{5}{2}\]